Problem 3 The USDA collects data on commodity prices and consumption. Use their data to explain per-capita chicken consumption as a function of the price of chicken. (Note that economic theory suggests that an increase in a commodity price is generally associated with a decrease in demand.) 0 per capita chicken consumption PCHICK consumer price index for chicken We estimated a linear regression model using annual data for 1950-2001 (N 52). Dependent variable: coefficlent std. error const РСНСК -0.1 0.15 R-squared: 0.854880 a, write the eason for the leastsquares regression line: У | b,The nul hypothesis s: β c·Calculate the t-rato- d. The crtcalvale,evaluated at α-0.05 for the dstributio(wth N-2 d)is e. Can we reject the null hypothesis? lpick one)
roblem 3 The USDA collects data on commodity prices and consumption. Use their data to explain per-capita chicken consumption as a function of the price of chicken. (Note that economic theory suggests that an increase in a commodity price is generally associated with a decrease in demand.) Q = per capita chicken consumption CHICKconsumer price index for chicken Collase A We estimated a linear regression model using annual data for 1950-2001 (N-52 Dependent variable: coefficient std. error const PCHICK -0.1 0.15 R-squared: 0.854880 a. Write the equation for the least-squares regression line: y= b, The null hypothesis is: β lock one)Bo c. Calculate the t-ratio d. The critical value, evaluated ata0.05, for the dstribution (with N-2 dr) is e. Can we reject the null hypotresis?(pick one)
Answer #1

a)

y^=intercept+slope*X

so, y^ = 5 - 0.1x

b)null hypothesis: ß = 0

c) t-ratio=estimated slope/SE = -0.1/0.15 = -0.666667

d) critical value

tα/2,df = t0.025,50= 2.00856

e)

| t-stat | < t-critical, so do not reject Ho

answer is NO